# Efficient Simulation of Open Quantum Systems on NISQ Devices

### v1

@@ III.3 Damped Harmonic Oscillator with Error Mitigation
[…]
@@
One simple method for mitigating device noise is to model the noise as some CPTP channel $\mathcal{E}(\rho)$ and attempt to partially mitigate device noise by inverting the channel. The inverted channel is applied to the noisy density matrix $\rho$ to obtain a mitigated density matrix $\rho_{\text{mit}}$. As we have seen in Section III.1, one such model is given by a Pauli channel. The evolution of a density matrix $\rho$ for a constant time period under a Pauli channel with Lindblad operators $L_{n}$ of the form (30) can be always written in the simplified form
@@
for all density matrices $\rho$. This means that one can correct for device-induced depolarizing errors that occur before and during the simulation by applying an inverse depolarization channel to the measured density matrix after simulation is complete. Moreover, inverting $\mathcal{E}_{\text{QDC}}$ can be carried out quite efficiently, requiring only a re-scaling of the diagonal and off-diagonal components of $\rho$ to obtain the mitigated density matrix $\rho_{\text{mit}}$. Although the Pauli channel model (38) is a much more flexible noise model, it does not exhibit the same commutativity as in (41) except in the case when $\mathcal{E}_{\text{sim}}$ is also a Pauli noise model. For more information on the process of fitting QDCs and Pauli channels to measurements on quantum hardware, we refer the reader to Appendix B. In the remainder of this section, we discuss the results of two experiments that apply the error mitigation techniques discussed above to the damped harmonic oscillator system.
@@ III.3.1 Noisy Oscillating State
To mitigate the effect of noise decoherence on the Harmony device, we fit a Pauli channel model and a QDC model to the initial state and select the respective parameters $\varepsilon_{i}$ and $\lambda$ such that the fidelity of the density matrix reconstructed on the $t=0$ density matrix is maximized. The results of the estimated Pauli channel and QDC parameters are shown in Figure 6(a) and 6(b) respectively. Since the QDC satisfies the property (40), a estimate of the parameter $\lambda$ that is sufficient to at least correct for decoherence during state preparation and measurement can be obtained by selecting the smallest value of $\lambda$ that maximizes the fidelity of the initial density matrix. As shown in Figure 6(b), estimating $\lambda$ in this manner approximately maximizes the average fidelity of the mitigated density matrices across the entire trajectory. In Figure 6(c) and 6(d), we plot the fidelity and von Neumann entropy of the density matrices estimated on the noisy device. In these figures, the gap between the noisy device results and a simulated ideal device are shaded in red, illustrating the effect of device noise decoherence. We observe in Figure 6(d) that both the Pauli channel and QDC error models decrease the density matrix entropy to closer resemble that of an ideal device. These methods also increase the average fidelity of the reconstructed density as seen in Figure 6(d), especially in the regime where the entropy of the density matrix is low. On the other hand, in the high-entropy regime (where $t\approx 0.5$) both error mitigation techniques reduce the entropy below that of the exact solution, resulting in a lower density matrix fidelity than the raw unmitigated data. This is illustrative of a key trade-off that exists between the fidelity of the mitigated density matrix in the high-entropy regime and low-entropy regimes: as the degree of mitigation is increased (i.e. the noise channel parameters $\varepsilon_{i}$ and $\lambda$ are increased), the fidelity in the high-entropy regime increases at the cost of some fidelity of the lower-entropy states. We remark that the exact nature of this trade-off, however, is likely to be specific to both the system being simulated and the quantum hardware it is simulated on.
@@
In Figure 11, we plot the trajectory of the position density and momentum density of the harmonic oscillator system interpolated as function of time. For each discrete time sample, the position and momentum densities were computed using (37) and interpolated between time steps with Gaussian smoothing. For comparison, the exact solution to the Lindblad equation is plotted in the last row of Figure 11. Although the unmitigated results appear to be degraded with noise, we see that much of this noise is suppressed after the inverse QDC error mitigation is applied. Nonetheless, the effect of noise is still prominent in the high-entropy regime at $t\approx 1.0$.
@@ III.3.2 Noisy Cat State
To provide a more concrete overview of how this can be incorporated into the error mitigation methods presented in the previous section, we present results obtained from a noisy emulation of the Quantinuum H1-1 device for the initial state (45) with $\alpha=1.2$ and $\gamma_{1}=0.6$. We simulate this system using the Kraus series method for nine uniform time steps from $t=0.0$ to $t=2.0$. In Figure 10 we plot the fidelity and entropy of the cat state trajectory estimated from the noisy device emulation and consider the effects of applying the twirling operation (46) and subsequent QDC inverse error mitigation. From these results, we note that twirling alone provides only a modest increase in overall state fidelity; however, when it is combined with the QDC error mitigation scheme, it provides a significant boost to the overall state fidelity, especially for the initial prepared state at $t=0$.
@@
Figure 10: Fidelity (a) and von Neumann Entropy (b) of the density matrix trajectory obtained through a noise model simulation of the Quantinuum H1-1 device. Results are shown for an ideal noiseless device and the simulated noisy device with no mitigation, with twirling, and with combined twirling and error mitigation. Error mitigation was applied using parity twirling and QDC inversion with a value of $\lambda_{\text{est}}=0.48$, which was obtained by fitting a QDC to the first state. The ideal trajectory of the system was obtained through a noiseless simulator with the same number of measurements as the noisy device.
@@
In Figure 12 we plot the estimated Wigner quasi-probability distributions of the emulated noisy device trajectory, and we compare these distributions to those of an ideal device. In Figure 10), we observed that even the initial state at $t=0$ was reproduced with low fidelity. From inspecting the Wigner distributions, we observe that this low fidelity partially manifests as broken parity symmetry. When parity twirling is applied to a harmonic oscillator state, the Wigner distribution of that state is averaged with its own $180^{\circ}$ rotation. It follows that an oscillator state exhibits parity symmetry if its Wigner distribution is invariant under $180^{\circ}$ rotations. We observe that this is the case for both the mitigated and ideal Wigner distributions, however the mitigated results at around $t=0$ exhibit distributions that appear to be more consistent with a single excited quantum state (i.e. $\rho\approx\ket{1}\bra{1}$ |) than with the approximate cat state (45). The Wigner distributions of cat states are characterized by the interference pattern of alternating positive and negative regions in phase space which separate two regions of high quasi-probability mass, which are not discernible in the mitigated results for $t<1.0$. Although we observe these deviations when the system is in the quantum mechanical regime at $t<1.0$, we also see that in the classical statistical regime at times $t>1.0$, there is strong agreement between the mitigated results and the ideal results. This suggests that parity twirling and subsequent error mitigation is very effective when applied to noisy simulations of systems with more classical character, but may not be able to restore all of the important quantum-mechanical effects in systems with a strong quantum character.
@@ Appendix A Mapping the Kraus Series to Quantum Circuits
First, we consider the problem of realizing the effective evolution operator $T(t)$. As discussed in Section II.3.1, the commutation relations (i) and (ii) ensure that the effective Hamiltonian $H_{\text{eff}}$ (given in (9)) can be diagonalized in the same basis as the system Hamiltonian $H$. It follows that $T(t)$ can then be diagonalized in the form
@@
where $W(t)$ is diagonal unitary and $\Lambda(t)$ is diagonal positive semi-definite with time-dependent entries of the form $W(t)_{ii}=e^{-i\theta_{i}t}$ and $\Lambda(t)_{ii}=e^{-\beta_{i}t}$, where $\theta_{i}\in[0,2\pi)$ and $\beta_{i}\geq 0$ correspond to the real and imaginary and components of $-itE_{\text{eff}}/\hbar$, where $E_{\text{eff}}$ (given in (10)) is an eigenvalue of $H_{\text{eff}}$. While one could in theory apply the diagonalizing unitary $U$ in the quantum circuit, this is often impractical on NISQ devices, since one might as well represent the entire system in this diagonal basis and avoid the expense of applying $U$ altogether.
@@
Methods of efficiently implementing diagonal unitaries $W(t)$ have been discussed in the literature (e.g., see [54]). Among these methods, there exists a general trade-off between classical pre-processing complexity and quantum gate complexity, depending on whether the diagonal elements are encoded with parameterized rotation gates via a standard binary or a Gray code encoding. In a binary encoding, the elements $W_{ii}(t)$ of a $d$-dimensional unitary are encoded through a sequence of $d$ multi-qubit controlled phase gates, where the control qubits for the $n$th phase gate are the $1$ bits in the binary representation of $n$. In this encoding, the phase angles corresponding to the entries $W_{ii}(t)$ be computed efficiently in time $O(d\log(d))$ and can be executed with $O(d\log(d))$ quantum gates [21]. In a Gray code encoding, controlled rotation gates are applied to single qubits such that a sequence of parity-controlled $R_{z}$ gates are applied to recursive sub-blocks of $W$ [54]. An important benefit of the Gray code encoding is that it requires roughly $d/2$ CNOT gates when $d$ is a power of $2$; however, the mapping of the entries $W_{i}i(t)$ to rotation gate parameters requires solving a $d^{2}\times d^{2}$ linear system, which has a classical complexity between $O(d^{2})$ and $O(d^{3})$ that dominates the quantum simulation time for large systems. For this reason, we consider the standard binary encoding more favorable when simulating the trajectory of large systems over many $t$ values, and the Gray code encoding more favorable for small systems with few $t$ values. In the results presented in Section III. The non-unitary diagonal matrix $\Lambda(t)$ can also be mapped to quantum gates through methods similar to those for the unitary $W(t)$; however instead of applying phase or $R_{z}$ gates, controlled $R_{y}$ gates are instead applied on an ancilla qubit with the system qubit as a control. This implements a unitary block encoding of the real diagonal operator $\Lambda(t)$. Both the standard binary encoding and the Gray code encoding of this non-unitary operator exhibit similar asymptotic quantum gate complexity and classical parameter mapping complexity as the diagonal opertor $W(t)$. For additional details regarding the trade-off, and for circuit diagrams implementing $\Lambda(t)$ and $W(t)$ for simple systems (such as the damped quantum harmonic oscillator), we refer the reader to reference [21].
@@ Appendix B Pauli Noise Models on NISQ Hardware
In this appendix, we consider Pauli noise models, and how they can be applied to density matrix simulations of open quantum systems on NISQ hardware. In Section III.3, we introduced the general Pauli channel noise model (38), which represents the Markovian evolution of a system under a continuous-time Pauli channel for some unknown time. We also considered a special case of the Pauli channel model, called the quantum depolarizing channel (QDC), given by (40). Although it is a relatively naive noise model, a QDC can be applied to model depolarization of quantum states, which is a type of decoherence that commutes with all other types of decoherence described by a CPTP map. Specifically, we claimed that the QDC satisfies the commutativity property (41), where $\mathcal{E}_{\text{sim}}$ is the CPTP map being simulated on the quantum device. This can be shown by expanding $\mathcal{E}_{\text{sim}}$ as a sum of Kraus operators $K_{i}$, so that by the unitality of the Kraus operators we obtain
@@
as desired. In a similar manner, it can also be shown that any two Pauli channels commute. Letting $P_{0}\equiv I$ for convenience of notation, we see that
@@
as desired. This result should come as no surprise, because we have already shown that the Lindblad operators $L_{n}$ of a continuous-time Pauli channel form a commutative group, hence the CPTP maps they generate should also commute.
@@
$A_{\mathcal{E}}=\sum_{i}(P_{i}\otimes\overline{P_{i}})$ (52)
@@
In this paper, we maximize (55) to find the optimal estimated value of $\lambda$ for the QDC as shown in Figure 6(b). For the general Pauli channel, however, we use the non-negative least squares approach that solves (50) for the optimal coefficients $\varepsilon_{i}$, as shown in Figure 6.

### v2

@@ III.3 Damped Harmonic Oscillator with Error Mitigation
So far, we have simulated relatively small systems on NISQ hardware, where the effect of noise is on the same order of magnitude as the sampling error. However, when simulating much larger systems on NISQ devices, the effect of noise is more problematic and must be combated through techniques such as hardware-native gate optimization and error mitigation [46, 47]. In this section, we focus on error-mitigation techniques, as circuit optimization is more hardware specific. In particular, we propose hardware-agnostic error mitigation techniques that can be applied to improve the accuracy of the density matrix trajectories of bosonic systems simulated using the Kraus series method. We apply these techniques to a damped harmonic oscillator system with Hamiltonian Eq. (16) and Lindblad operator $L_{1}=\hat{a}$ with damping parameter $\gamma_{1}$. We restrict the representation of $\hat{a}$ to incorporate occupied eigenstates up to $n=3$ in a dense two-qubit binary encoding. This allows for a Kraus representation of the form in Eq. (6) with $f(t)=(1-e^{-\gamma_{1}t}))/\gamma_{1}$.
@@
One simple method for mitigating device noise is to model the noise as some CPTP channel $\mathcal{E}(\rho)$ and attempt to partially mitigate device noise by inverting the channel. The inverted channel is applied to the noisy density matrix $\rho$ to obtain a mitigated density matrix $\rho_{\text{mit}}$. As we have seen in Section III.1, one such model is given by a Pauli channel. The evolution of a density matrix $\rho$ for a constant time period under a Pauli channel with Lindblad operators $L_{n}$ of the form in Eq. (30) can be always written in the simplified form
@@
for all density matrices $\rho$. This means that one can correct for device-induced depolarizing errors that occur before and during the simulation by applying an inverse depolarization channel to the measured density matrix after simulation is complete. Moreover, inverting $\mathcal{E}_{\text{QDC}}$ can be carried out quite efficiently, requiring only a re-scaling of the diagonal and off-diagonal components of $\rho$ to obtain the mitigated density matrix $\rho_{\text{mit}}$. Although the Pauli channel model in Eq. (38) is a much more flexible noise model, it does not exhibit the same commutativity as in Eq. (41) except in the case when $\mathcal{E}_{\text{sim}}$ is also a Pauli noise model. For more information on the process of fitting QDCs and Pauli channels to measurements on quantum hardware, we refer the reader to Appendix B. In the remainder of this section, we discuss the results of two experiments that apply the error mitigation techniques discussed above to the damped harmonic oscillator system.
@@ III.3.1 Noisy Oscillating State
To mitigate the effect of noise decoherence on the Harmony device, we fit a Pauli channel model and a QDC model to the initial state and select the respective parameters $\varepsilon_{i}$ and $\lambda$ such that the fidelity of the density matrix reconstructed on the $t=0$ density matrix is maximized. The results of the estimated Pauli channel and QDC parameters are shown in Figure 6(a) and 6(b) respectively. Since the QDC satisfies Eq. (40), an estimate of the parameter $\lambda$ that is sufficient to at least correct for decoherence during state preparation and measurement can be obtained by selecting the smallest value of $\lambda$ that maximizes the fidelity of the initial density matrix. As shown in Figure 6(b), estimating $\lambda$ in this manner approximately maximizes the average fidelity of the mitigated density matrices across the entire trajectory. In Figure 6(c) and 6(d), we plot the fidelity and von Neumann entropy of the density matrices estimated on the noisy device. In these figures, the gap between the noisy device results and a simulated ideal device are shaded in red, illustrating the effect of device noise decoherence. We observe in Figure 6(d) that both the Pauli channel and QDC error models decrease the density matrix entropy to closer resemble that of an ideal device. These methods also increase the average fidelity of the reconstructed density as seen in Figure 6(d), especially in the regime where the entropy of the density matrix is low. On the other hand, in the high-entropy regime (where $t\approx 0.5$) both error mitigation techniques reduce the entropy below that of the exact solution, resulting in a lower density matrix fidelity than the raw unmitigated data. This is illustrative of a key trade-off that exists between the fidelity of the mitigated density matrix in the high-entropy regime and low-entropy regimes: as the degree of mitigation is increased (i.e. the noise channel parameters $\varepsilon_{i}$ and $\lambda$ are increased), the fidelity in the high-entropy regime increases at the cost of some fidelity of the lower-entropy states. We remark that the exact nature of this trade-off, however, is likely to be specific to both the system being simulated and the quantum hardware it is simulated on.
@@
In Figure 11, we plot the trajectory of the position density and momentum density of the harmonic oscillator system interpolated as function of time. For each discrete time sample, the position and momentum densities were computed using Eq. (37) and interpolated between time steps with Gaussian smoothing. For comparison, the exact solution to the Lindblad equation is plotted in the last row of Figure 11. Although the unmitigated results appear to be degraded with noise, we see that much of this noise is suppressed after the inverse QDC error mitigation is applied. Nonetheless, the effect of noise is still prominent in the high-entropy regime at $t\approx 1.0$.
@@ III.3.2 Noisy Cat State
To provide a more concrete overview of how this can be incorporated into the error mitigation methods presented in the previous section, we present results obtained from a noisy emulation of the Quantinuum H1-1 device for the initial state in Eq. (45) with $\alpha=1.2$ and $\gamma_{1}=0.6$. We simulate this system using the Kraus series method for nine uniform time steps from $t=0.0$ to $t=2.0$. In Figure 10 we plot the fidelity and entropy of the cat state trajectory estimated from the noisy device emulation and consider the effects of applying the twirling operation in Eq. (46) and subsequent QDC inverse error mitigation. From these results, we note that twirling alone provides only a modest increase in overall state fidelity; however, when it is combined with the QDC error mitigation scheme, it provides a significant boost to the overall state fidelity, especially for the initial prepared state at $t=0$.
@@
Figure 10: Fidelity (a) and von Neumann Entropy (b) of the density matrix trajectory obtained through a noise model emulation of the Quantinuum H1-1 device. Results are shown for an ideal noiseless device and the simulated noisy device with no mitigation, with twirling, and with combined twirling and error mitigation. Error mitigation was applied using parity twirling and QDC inversion with a value of $\lambda_{\text{est}}=0.48$, which was obtained by fitting a QDC to the first state. The ideal trajectory of the system was obtained through a noiseless simulator with the same number of measurements as the noisy device.
@@
In Figure 12 we plot the estimated Wigner quasi-probability distributions of the emulated noisy device trajectory, and we compare these distributions to those of an ideal device. In Figure 10), we observed that even the initial state at $t=0$ was reproduced with low fidelity. From inspecting the Wigner distributions, we observe that this low fidelity partially manifests as broken parity symmetry. When parity twirling is applied to a harmonic oscillator state, the Wigner distribution of that state is averaged with its own $180^{\circ}$ rotation. It follows that an oscillator state exhibits parity symmetry if its Wigner distribution is invariant under $180^{\circ}$ rotations. We observe that this is the case for both the mitigated and ideal Wigner distributions, however the mitigated results at around $t=0$ exhibit distributions that appear to be more consistent with a single excited quantum state (i.e. $\rho\approx\ket{1}\bra{1}$ |) than with the approximate cat state in Eq. (45). The Wigner distributions of cat states are characterized by the interference pattern of alternating positive and negative regions in phase space which separate two regions of high quasi-probability mass, which are not discernible in the mitigated results for $t<1.0$. Although we observe these deviations when the system is in the quantum mechanical regime at $t<1.0$, we also see that in the classical statistical regime at times $t>1.0$, there is strong agreement between the mitigated results and the ideal results. This suggests that parity twirling and subsequent error mitigation is very effective when applied to noisy simulations of systems with more classical character, but may not be able to restore all of the important quantum-mechanical effects in systems with a strong quantum character.
@@ IV Conclusion
Data Availability Statement
The code, data, and demo notebooks necessary to reproduce these results can be found in the associated Github repository [54].
@@ Acknowledgements.
Author Contributions
C.B. and E.P.B conceptualized the work. C.B. developed the theory and code. C.B., N.B., and G.S. conducted the quantum hardware experiments. All authors contributed to the writing and review of the manuscript.
@@ Appendix A Mapping the Kraus Series to Quantum Circuits
First, we consider the problem of realizing the effective evolution operator $T(t)$. As discussed in Section II.3.1, the commutation relations (i) and (ii) ensure that the effective Hamiltonian $H_{\text{eff}}$ (given in Eq. (9)) can be diagonalized in the same basis as the system Hamiltonian $H$. It follows that $T(t)$ can then be diagonalized in the form
@@
where $W(t)$ is diagonal unitary and $\Lambda(t)$ is diagonal positive semi-definite with time-dependent entries of the form $W_{ii}(t)=e^{-i\theta_{i}t}$ and $\Lambda_{ii}(t)=e^{-\beta_{i}t}$, where $\theta_{i}\in[0,2\pi)$ and $\beta_{i}\geq 0$ correspond to the imaginary and real components of $-itE_{\text{eff}}/\hbar$, where $E_{\text{eff}}$ (given in Eq. (10)) is an eigenvalue of $H_{\text{eff}}$. While one could in theory apply the diagonalizing unitary $U$ in the quantum circuit, this is often impractical on NISQ devices, since one might as well represent the entire system in this diagonal basis and avoid the expense of applying $U$ altogether.
@@
Methods of efficiently implementing diagonal unitaries $W(t)$ have been discussed in the literature (e.g., see [55]). Among these methods, there exists a general trade-off between classical pre-processing complexity and quantum gate complexity, depending on whether the diagonal elements are encoded with parameterized rotation gates via a standard binary or a Gray code encoding. In a binary encoding, the elements $W_{ii}(t)$ of a $d$-dimensional unitary are encoded through a sequence of $d$ multi-qubit controlled phase gates, where the control qubits for the $n$th phase gate are the $1$ bits in the binary representation of $n$. In this encoding, the phase angles corresponding to the entries $W_{ii}(t)$ can be computed efficiently in time $O(d\log(d))$ and can be executed with $O(d\log(d))$ quantum gates [21]. In a Gray code encoding, controlled rotation gates are applied to single qubits such that a sequence of parity-controlled $R_{z}$ gates are applied to recursive sub-blocks of $W$ [55]. An important benefit of the Gray code encoding is that it requires roughly $d/2$ CNOT gates when $d$ is a power of $2$; however, the mapping of the entries $W_{i}i(t)$ to rotation gate parameters requires solving a $d^{2}\times d^{2}$ linear system, which has a classical complexity between $O(d^{2})$ and $O(d^{3})$ that dominates the quantum simulation time for large systems. For this reason, we consider the standard binary encoding more favorable when simulating the trajectory of large systems over many $t$ values, and the Gray code encoding more favorable for small systems with few $t$ values. In the results presented in Section III. The non-unitary diagonal matrix $\Lambda(t)$ can also be mapped to quantum gates through methods similar to those for the unitary $W(t)$; however instead of applying phase or $R_{z}$ gates, controlled $R_{y}$ gates are instead applied on an ancilla qubit with the system qubit as a control. This implements a unitary block encoding of the real diagonal operator $\Lambda(t)$. Both the standard binary encoding and the Gray code encoding of this non-unitary operator exhibit similar asymptotic quantum gate complexity and classical parameter mapping complexity as the diagonal opertor $W(t)$. For additional details regarding the trade-off, and for circuit diagrams implementing $\Lambda(t)$ and $W(t)$ for simple systems (such as the damped quantum harmonic oscillator), we refer the reader to reference [21].
@@ Appendix B Pauli Noise Models on NISQ Hardware
In this appendix, we consider Pauli noise models, and how they can be applied to density matrix simulations of open quantum systems on NISQ hardware. In Section III.3, we introduced the general Pauli channel noise model Eq. (38), which represents the Markovian evolution of a system under a continuous-time Pauli channel for some unknown time. We also considered a special case of the Pauli channel model, called the quantum depolarizing channel (QDC), given by Eq. (40). Although it is a relatively naive noise model, a QDC can be applied to model depolarization of quantum states, which is a type of decoherence that commutes with all other types of decoherence described by a CPTP map. Specifically, we claimed that the QDC satisfies the commutativity property Eq. (41), where $\mathcal{E}_{\text{sim}}$ is the CPTP map being simulated on the quantum device. This can be shown by expanding $\mathcal{E}_{\text{sim}}$ as a sum of Kraus operators $K_{i}$, so that by the unitality of the Kraus operators we obtain
@@
as claimed. In a similar manner, it can also be shown that any two Pauli channels commute. Letting $P_{0}\equiv I$ for convenience of notation, we see that
@@
as claimed. This result should come as no surprise, because we have already shown that the Lindblad operators $L_{n}$ of a continuous-time Pauli channel form a commutative group, hence the CPTP maps they generate should also commute.
@@
$A_{\mathcal{E}}=\sum_{i}\varepsilon_{i}(P_{i}\otimes\overline{P_{i}})$ (52)
@@
In this paper, we maximize Eq. (55) to find the optimal estimated value of $\lambda$ for the QDC as shown in Figure 6(b). For the general Pauli channel, however, we use the non-negative least squares approach that solves Eq. (50) for the optimal coefficients $\varepsilon_{i}$, as shown in Figure 6.